OpenStudy (anonymous):

With x-intercept =-3 and y =4 Find the equationof the line in slope intercept form

OpenStudy (mertsj):

If the x-intercept is -3, what ordered pair is on the line?

OpenStudy (anonymous):

3,0

OpenStudy (mertsj):

(-3,0)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

The first thing is, you'll have to get the slope:\[m = \frac{ y _{2} - y _{1} }{ x _{2} - x _{1} }\]where you have 2 points: (-3, 0) and (0, 4) which correspond to your intercepts.

OpenStudy (mertsj):

And if the y intercept is 4, what ordered pair is on the line?

OpenStudy (anonymous):

Next, you'll first have to use the point-slope equation of a line:\[y - y _{1} = m(x - x _{1})\]and you can use either of the two points for this. Last, we can work on converting that equation to slope-intercept form: y = mx + b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, your slope is:\[m = \frac{ 4 - 0 }{ 0 - (-3) } = 4/3\]

OpenStudy (anonymous):

well tht was easy

OpenStudy (anonymous):

That first equation point-slope, the last is slope-intercept. And that's it!

OpenStudy (anonymous):

ok what about parallel to the line 4x +y=6, and through the point (-3,5)

OpenStudy (anonymous):

Wait, hold on, I have to re-do something.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The point-slope form is: y - 4 = (4/3)(x - 0) So, the slope-intercept form is: y - 4 = (4/3)x -> y = (4/3)x + 4

OpenStudy (anonymous):

I had a typo before.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok, I see your next one, and I'll do it here, but the next ones should be in separate threads. We have to get the slope and that is gotten by putting equation in slope-intercept form: y = -4x + 6 where it is in the form: y = mx + b So, slope is -4 That's step #1.

OpenStudy (anonymous):

Now we use point-slope again: y - 5 = (-4)[x - (-3)] Now, we put that into slope-intercept: y - 5 = -4x - 12 y = -4x - 7 And that's done. Ok, "new thread time" unless you have a question on this last one we just did.

OpenStudy (anonymous):

ok new theard